OpenStudy (anonymous):

The formula for determining interest compounded monthly is A = P(1 + )12t, where A represents the amount invested after t years, P the principal invested, and r the interest rate. Jimmy invests $1,000 at an interest rate of 10% for 3 years, while Jenny invests $1,000 at an interest rate of 5% for 6 years. Part 1: Determine the amount of return gained by Jimmy and Jenny. Part 2: Summarize your results from Part 1, including how you arrived at your answer.

OpenStudy (helder_edwin):

if you principal is P, after one year you have \[ \Large P+rP=(1+r)P \] after the second year you have \[ \Large (1+r)P+r(1+r)P=(1+r)(1+r)P=(1+r)^2P \] so in general you have, after n years \[ \Large (1+r)^nP \] do u understand?